OpenStudy (cakecup2018):

I NEED HELP !!! WITH RELATIVE ERROR, ABSOLUTE ERROR, AND PERCENT ERROR, ! WILL GIVE A METAL!!!

OpenStudy (whpalmer4):

What's the question, specifically?

OpenStudy (cakecup2018):

Well its not specific questions i just need the whole subject, but i can give you an example

OpenStudy (whpalmer4):

okay, fire away

OpenStudy (cakecup2018):

Bob was told to measure 45 cm of copper wire to use in an experiment. Since his ruler only measured to 48 cm he used this amount, and his experiment was a failure. How much error did Bob introduce into his experiment?

OpenStudy (cakecup2018):

give the absolute error, relative error, and percent error..

OpenStudy (cakecup2018):

i have my exam tomorrow so i need help with a couple other things too

OpenStudy (whpalmer4):

Well, absolute error is the difference between your measurement and the correct (or accepted) amount. It is always a positive number, so it is the absolute value of your measurement minus the correct value. \[\text{absolute error} = |\text{measurement} - \text{correct value}|\] Relative error is the fraction of absolute error divided by correct value. \[\text{relative error} = \frac{\text{absolute error}}{\text{correct value}} = \frac{|\text{measurement} - \text{correct value}|}{\text{correct value}}\] Percent error is simply relative error expressed as a percentage: \[\text{percent error} = \text{relative error} * 100\%\]or\[\text{percent error} = \frac{|\text{measurement} - \text{correct value}|}{\text{correct value}}*100\%\]

OpenStudy (cakecup2018):

so can we do it the example together?

OpenStudy (cakecup2018):

r u there @whpalmer4 ??

OpenStudy (whpalmer4):

okay, use what I just said and come up with answers. I'll tell you if I agree or not.

OpenStudy (cakecup2018):

0.0625

OpenStudy (whpalmer4):

Okay, the problem asks for 3 different results. You just gave me 1 number. I would mark that wrong, even if it turns out to be one of the 3 numbers, because you didn't tell me which one it is.

OpenStudy (whpalmer4):

Examples of good and bad answers: Bad: 0.0625 Good: The relative error is 0.0625.